# LayerCollapse: Adaptive compression of neural networks

### v1

@@ Abstract
Handling the ever-increasing scale of contemporary deep learning and transformer-based models poses a significant challenge. Although great strides have been made in optimizing model compression techniques such as model architecture search and knowledge distillation, the availability of data and computational resources remains a considerable hurdle for these optimizations. This paper introduces LayerCollapse, a novel alternative adaptive model compression methodology. LayerCollapse works by eliminating non-linearities within the network and collapsing two consecutive fully connected layers into a single linear transformation. This approach simultaneously reduces both the number of layers and the parameter count, thereby enhancing model efficiency. We also introduce a compression aware regularizer, which compresses the model in alignment with the dataset quality and model expressiveness, consequently reducing overfitting across tasks. Our results demonstrate LayerCollapse’s effective compression and regularization capabilities in multiple fine-grained classification benchmarks, achieving up to 74% post training compression with minimal accuracy loss. We compare this method with knowledge distillation on the same target network, showcasing a five-fold increase in computational efficiency and 8% improvement in overall accuracy on the ImageNet dataset.
@@ 1 Introduction
The expanding scope of Deep Learning (DL) applications has led to an escalating demand for more accurate models, resulting in an ongoing rise in their complexity. This surge in complexity, however, presents a challenge: it increases the execution costs, making it difficult to deploy these models in real-world scenarios on standard hardware. For example, transformer-based deep models while improving accuracy, present new computational challenges. Fortunately, modern trained neural networks are shown to include excess redundancies that can be removed without losing accuracy. This potential for reduction has catalyzed research in model compression strategies such as pruning [24][19], quantization [15][18][29][20], and model distillation[14], focusing on the effective detection and elimination of these superfluous components. Despite the amazing progress in the field of model compression, these optimizations demand the availability of data and high computational resources which further hinders their practicability.
Multi-Layer Perceptrons (MLPs) [11] is a key architectural component of modern deep neural networks. MLPs are regarded by many as the first contribution to deep learning, first introduced in 1965 by Ivakhnenko [17][32]. In its simplest form, a $K$-layer MLP can be described as $K$ linear transformations separated by some non-linear operations. Equivalently, an MLP can also be described as a sequence of fully connected (a.k.a., dense), layers separated by non-linear layers. Often, MLP layers serve as the final decision-making components, such as classification heads. MLPs are central to the success of deep learning models in diverse architectures including Convolutional Neural Networks [33], Vision Transformers [7], and MLP-Mixers [35].
Because of the dense nature of MLPs, a significant portion of a model’s computational overhead and storage requirements can be attributed to the MLP layers. For instance, MLPs constitute over 60% of the model’s total parameters for the popular models discussed in [7][35][33].
Table 1: Percentage of model parameters and MACs reductions attainable by collapsing their MLP layers
@@
MLPs contain non-linear activations, e.g., ReLU layers, which separate sequences of linear transformations. If the non-linear layers were absent, the consecutive linear layers could be condensed into a single layer.To showcase the immense potential of computation and storage reductions when linearizing MLPs, we calculate this reduction in the architectures presented in [7][35][33]. Table 1 demonstrates the reduction percentage in the number of parameters and MAC operations when we replace the MLP’s with a linear transform in relevant vision architectures.
The core thesis of our work revolves around non-linear activations. We argue that the expressiveness of an MLP is primarily governed by its non-linear components. By modulating these non-linear characteristics, we can orchestrate a continuum of model expressiveness, ranging from highly complex to near-linear transformations. Figure 1 demonstrates this idea using a simple regression task and an MLP with 2 layers. As illustrated, having complex models with lack of a sizeable dataset will lead to overfitting. By modulating the non-linear activation layer one can regulate model’s expressiveness.
@@
Figure 1: Illustration of regression performance on a two-layer neural network, demonstrating the impact of varying levels of ReLU activation linearity on model overfitting and underfitting.
Classically, regularization strategies such as the $\ell_{1}$ and $\ell_{2}$ norms, have been employed to restrict model expressiveness, ensuring robust generalization and preventing overfitting. These techniques also contribute to compression-aware training. In similar veins, Lemaire et. al. describe a budget aware regularizer [23] and Nguyen et. al. introduce a regularizer for quantization aware training [29]. Extending this principle, we introduce a compression aware regularizer by using the nature of ReLU activations. This regularizer works by shifting the slope of Relu towards linearity. This not only facilitates the fusion of successive linear layers, conserving computational resources and storage, but also can lead to better model generalizability.
@@
This paper presents several key contributions to the field of neural network compression and regularization:
  • Introduction of a novel post-training technique, called LayerCollapse, that combines two fully connected layers into one, reducing model size and complexity.
  • Development of a novel compression aware regularization loss with the objective of promoting shallower structures during training.
  • Proof-of-concept experimentation on a diverse set of datasets benchmarks, in vision and NLP domaains, showing post training compression up to 74%. Furthermore we show improvement in generalizability when training with proposed compression aware regularizer while decreasing size upto 50%. Lastly we compare the computational performance of knowledge distillation and LayerCollapse showing an 8% final result improvement while using 80% less computational resources.
@@
  • Providing an open-source library that enables easy implementation of LayerCollapse, compatible with PyTorch. [^1] https://github.com/sohail2810/LayerCollapse
2 Related work
2.1 Compression
In the realm of deep learning, compression techniques aim to reduce computational demands while ensuring model precision. With the emergence of foundational models and the tapering of Moore’s law, the correct utilization of resources is a paramount concern. Predominantly, the methods employed for model compression are pruning [10], quantization [15][18], and distillation [14], each with its unique approach toward model optimization.
Pruning eliminates neurons whose impact on model performance are marginal or negligible. A big body of work explores diverse strategies to discern which neurons are expendable, including magnitude [24] and gradient [26][41] based methods. Another body of work looks into pruning aware training (PAT). In these methods a regularization term ensures structural sparsity when training [23][19][38][1]. This reduces the loss of accuracy when pruning the model post training.
Quantization constrains the parameter range and maps it into integers, thereby reducing the model’s computational and storage requirements. This methodology has been explored and implemented in various studies, providing numerous approaches toward effective quantization [9][29]. To minimize the loss induced by quantization it is standard practice to do quantization aware training (QAT). In short, we regularize the model weights during training to stay within a a close range of their quantized value [29].
Distillation[14], within another compression category, initially seeks an optimized architecture via Neural Architecture Search (NAS), a procedure notably resource-intensive in itself. The target model is trained using this new architecture, utilizing the dataset and the original model as a teacher. This process is sometimes called the student-teacher paradigm. It’s important to note that both NAS and distillation are resource-draining processes and require access to training data, which may not always be accessible.
In section 4, we compare LayerCollapse against distillation, highlighting its respective merits and applicability.
2.2 Regularization
In the field of deep learning, achieving the right balance between underfitting and overfitting is a pivotal challenge. Overfitting occurs when a model excessively learns from the noise in the training data, reducing its ability to generalize to new datasets. This issue is typically managed through the bias-variance trade-off, a principle that aids in creating models capable of effective generalization. A model with high variance is overly sensitive to small fluctuations in data, leading to unstable predictions. This trade-off highlights the necessity of maintaining an equilibrium between model complexity and the volume and integrity of the data. To counteract overfitting and enhance performance on test data, it’s crucial to implement explicit regularization strategies. These methods ensure the model’s robustness and consistent performance across both training and test datasets. Research by Yang et al. demonstrated that $\ell_{1}$ regularization can serve as a pruning-aware training method to sparsify the model, thereby boosting its pruning efficiency[40]. Ding et al. employed a specialized $\ell_{2}$ regularization to selectively eliminate less significant channels during training, aiming for subsequent pruning [4]. Additionally, Nguyen et al. introduced a regularization technique specifically for quantization-aware training [29]. Drawing inspiration from Santos et al.’s classification framework, we categorize model generalization techniques into three distinct categories: “Ensembling”, “Explicit Model Regularizers”, and “Dataset Design”[31].
Ensembling techniques often revolve around the integration of multiple models or sub-models to enhance performance and generalization. Within this domain, methods such as Dropout and its variants [34][5][8] play a pivotal role.
Dataset Design centers on refining the input data and labels to achieve model generalization. This includes data augmentation techniques such as transformative actions (random cropping, blurring, and rotation) seen in works like Bag of Tricks [13], FixRes [36], Cutout [3], CutMix [42], and AutoAugment [25].
Explicit regularizers modulate model configurations either through architectural adjustments or by penalizing certain weights to prevent overfitting. These methods follow the principle of Occam’s razor. They assume a prior knowledge that simpler models are more likely to be robust. This category encompasses classical regularization techniques like L0, L1, and L2 (weight decay) [27]. Our proposed method, falls into the same category. In section 3, we demonstrate how we use this principle in designing our regularization term. While classical norms influence weight magnitudes to mitigate overfitting, LayerCollapse uniquely stands out in its ability to adjust the model’s depth and structure.
@@ 3 Methodology
Figure 2: Overview of the LayerCollapse process: $(a)$ The compression aware the regularization loss shifts the PReLU towards linearity. $(b)$ The layer collapse operation involves eliminating a layer by replacing the hidden layer and adjusting the weights through the product of two matrices.
This section introduces LayerCollapse, an innovative compression method that adaptively reduces the model’s depth through regularization. This methodology aims to optimize the size of the model based on the quality of the dataset, resulting in enhanced generalization. Our proposed framework seamlessly integrates regularization and compression.
Unlike pruning [24] and quantization [15] which require the compressed model to be finetuned, LayerCollapse requires the finetuning prior to the compression phase.
As depicted in Figure 2, during a compression aware training or finetuning, we regularize a layer’s complexity by linearizing its activation function. We define Collapsing an MLP on a model that has shifted at least one of its non-linear layers into linearity. As depicted in Figure 2 Linearizing a non-linear activation within a model allows for the collapse of the two surrounding linear layers into one linear layer, effectively reducing the depth. The resulting model has fewer parameters and a smaller depth, requiring less resources for execution. For post training compression, one needs to run a few cycles of finetuning using the LayerCollapse regularizer to ensure the linearization loss is marginal.
We define a Collapsible MLP to be two consecutive fully connected layers utilizing a Parametric Rectified Linear Unit (PReLU) [12] in the middle. The Parametric Rectified Linear Unit (PReLU) can be expressed as:
@@
where, $W_{1}$ and $W_{2}$ are the weight matrices and $b_{1}$ and $b_{2}$ are the bias vectors for the first and second fully connected layers, respectively. $\alpha$ can be used to make the activation linear. We define $Y_{linear}$ as:
@@
We define the error as the squared difference of $Y_{\alpha}$ and $Y_{linear}$. In Appendix A, we show that $\forall\delta>0$
@@
where, $\sigma_{max}(W_{2}W_{1})$ is the largest singular value of $W_{2}W_{1}$ By choosing $\delta$ arbitrarily small, error is bounded proportional to $(1-\alpha)^{2}$ with probability one. We capitalize on this by incorporating $(1-\alpha)^{2}$ into our loss function. This loss term serves to align $Y_{\alpha}$ with $Y_{linear}$, compressing the model into a shallower model as demonstrated in Figure 2. In simpler terms, when $\alpha$ approaches one, substituting $Y_{linear}$ with $Y_{\alpha}$ with an error bounded by a constant multiplied by $(1-\alpha)^{2}$.
In practical applications, Dropout [34] and BatchNormalization [16] are frequently used between the layers of a network. Dropout works by temporarily deactivating random neurons during training, which helps spread the learning across different neurons to avoid over-reliance on any particular subset of them. It is important to note that Dropout is not active during model deployment, meaning it behaves as if it is not there. This characteristic means that Dropout doesn’t affect the layer collapsing process and can be left out when considering it.
@@
where, $\gamma$ and $\beta$ are learnable parameters, the $E[X]$ and $Var[X]$ are the mean and variance of the current batch, which we sometimes adjust with a method called Polyak averaging to keep the batch-to-batch changes smooth. Algorithm 1 outlines the specific steps to be taken for adjusting weights and biases when BatchNormalization is present.
To measure how well our compression method works, we use the concept of “compression gain”. This is calculated as the percentage reduction in the number of parameters of a layer, excluding bias and BatchNormalization parameters. The formula for compression gain is:
$Gain=1-\frac{n_{in}\times n_{out}}{n_{hidden}\times(n_{in}+n_{out})}$ (7)

### v3

@@ Abstract
Handling the ever-increasing scale of contemporary deep learning and transformer-based models poses a significant challenge. Overparameterized Transformer networks outperform prior art in Natural Language processing and Computer Vision. These models contain hundreds of millions of parameters, demanding significant computational resources and making them prone to overfitting on down stream tasks. In this work we present LayerCollapse , a novel structured pruning method to reduce the depth of fully connected layers. We propose an innovative regularizer that promotes shallow fully connected layers, compressing the model with minimal performance impact. This regularizer enables post-training compression without fine-tuning while preserving performance. LayerCollapse controls model expressiveness by regularizing the activation functions between fully connected layers, modulating them to linearity. A linear activation function collapses the rank of a transformation to the rank of the corresponding linear transformation, which demands less resources from the hardware. We demonstrate the effectiveness of LayerCollapse by showing its compression capabilities in sentimental analysis, text generation, and image classification benchmarks.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
The expanding scope of Deep Learning (DL) applications has led to an escalating demand for more accurate models, resulting in an ongoing rise in their complexity. This surge in complexity, however, presents a challenge: It increases the execution costs, in terms of memory and latency. This makes it difficult to deploy these models in real-world scenarios on standard hardware. Recent studies have shown that modern neural networks often contain unnecessary redundancies, which can be eliminated without affecting their performance. This discovery has spurred interest in model compression techniques such as unstructured pruning Li et al. (2016) Jiang et al. (2023), structured pruning Fan et al. (2019), quantization Hubara et al. (2017) Jacob et al. (2018) Nguyen et al. (2020) Krishnamoorthi (2018), and model distillation Hinton et al. (2015). These methods aim to identify and remove the redundant elements effectively. However, despite significant advances in model compression, implementing the pertinent optimizations typically requires additional steps such as fine-tuning. This can be resource-intensive, making the compression techniques less practical. Moreover some of these methods like unstructured pruning require support from hardware and software to take advantage of the compression Sundaram et al. (2023).
Table 1: Percentage of model parameters and MACs associated with MLP layers of streamline architectures.
@@
Multi-Layer Perceptrons (MLPs) Haykin (1994) are a key architectural component of modern deep neural networks. MLPs are regarded by many as the first contribution to deep learning, present in modern deep learning designs such as the transformer architecture. In its simplest form, a $K$-layer is defined as as $K$ linear transformations separated by non-linear operations. MLPs are central to the success of diverse deep learning architectures including encoder only transformer models Vaswani et al. (2017) Lan et al. (2019) Raffel et al. (2020), causal transformer models Radford et al. (2019) Touvron et al. (2023) Gunasekar et al. (2023), Vision Transformers Dosovitskiy et al. (2020), and MLP-Mixers Tolstikhin et al. (2021). Due to the dense nature of MLPs, a significant portion of a model’s computational overhead and storage requirements can be attributed to its MLP layers. For instance, MLPs constitute over 60% of the model’s total parameters for the popular architectures discussed in Dosovitskiy et al. (2020) Tolstikhin et al. (2021) Simonyan & Zisserman (2014). MLPs contain non-linear activations, e.g., ReLU, which separate sequences of linear transformations. If the non-linear layers were absent, the consecutive linear layers could be condensed into a single linear transformation. Table 1 showcases the portion of storage and computation requirements of MLP layers is the mentioned architectures.
@@
Figure 1: Illustration of regression performance on a two-layer neural network, demonstrating the impact of varying levels of ReLU activation linearity on model overfitting and underfitting. We define the percent of linearity to be the negative slope of the ReLU activation.
Figure 2: Overview of the LayerCollapse process: $(a)$ The regularization loss shifts the PReLU towards linearity. $(b)$ The layer collapse operation involves eliminating a layer by replacing the hidden layer and adjusting the weights through the product of two matrices.
In this work, we focus on the non-linear activations inside MLPs. We argue that the expressiveness of an MLP is primarily governed by its non-linear component. By modulating these non-linear characteristics, we can orchestrate a continuum of model expressiveness, ranging from highly complex to near-linear transformations. To gain intuition on the effect of non-linear activations control on model expressiveness, we use a simple 2-layer MLP for a simple regression task. The results are illustrated in figure 1. Here an $n\%$ linear model is defined as a model where the negative slope of the leaky ReLU is $n\%$ of one. So, the regular model is just a 2-layer MLP with ReLU activation and a $98\%$ linear model is a 2-layer MLP with a leaky ReLU activation with a negative slope of $0.98$. As we can see, different $n$ values create a continuum of model expressiveness. The regular model overfits to the noise within the data, however, the $80\%$ linear model fits the underlying distribution nicely while the $98\%$ linear model under-fits the data and cannot capture the complexity of the regression task.
A linear transformation can be performed using only a single matrix multiplication, while a $K$-layer MLP requires $K$ matrix multiplications. By adjusting the non-linear component towards linearity, we can regulate the model expressiveness while reducing the operational complexity of the model. We replace two matrix multiplications and a non-linear transform with only one smaller matrix multiplication when the modulated non-linearity is close to linear. In section 3, we will formally define close to linear and its effect on accuracy.
Controlling the model expressiveness using this technique will allow us to tailor the complexity of our model based on the task. Starting from the same architecture, some tasks have small datasets that require a less expressive model to prevent overfitting.
In this work, we developed a novel regularization technique to limit the model expressiveness to prevent overfitting, while reducing the computation cost of the model.
Common, regularization strategies such as Dropout Srivastava et al. (2014), $\ell_{1}$ and $\ell_{2}$ norms, have been employed to restrict model expressiveness, attempting robust generalization and preventing overfitting. These regularizers change the model weight structures such that post training compression like pruning and quantization pared with finetuning will yield efficient high performing models. Fan et al. introduced LayerDrop, a regularization technique to allow for post-training compression without fine-tuning Fan et al. (2019). LayerDrop employs a regularizer that controls the amount of bypass signals in each MLP layer within transformer architectures. Extending this idea, we introduce LayerCollapse regularization. The LayerCollapse regularizer works by modulating the negative slope of the ReLU non-linearity to one, effectively transforming it into a linear transformation. This modification not only facilitates the fusion of successive linear layers, conserving computational resources, but also can enhance model generalizability.
@@
This paper presents several key contributions to the field of neural network regularization and compression:
  • We introduce LayerCollapse , a novel regularization approach that enables effective one-shot compression after training.
  • We make a theoretical analysis providing bounds on the compression loss associated with LayerCollapse .
  • We propose a compression-aware regularization loss designed to encourage simpler model structures during the training and finetuning phase.
  • Through experiments across various datasets in both vision and NLP fields, we demonstrate that LayerCollapse achieves post-training compression rates of up to 74% for VGG models and 16% for transformer architectures. These experiments also highlight enhanced model generalization and reduced overfitting.
@@
  • We offer an open-source library for LayerCollapse , facilitating its easy integration with PyTorch frameworks. [^1] https://github.com/sohail2810/LayerCollapse
2 Background
Multi-Layer Perceptrons (MLPs) are foundational elements in modern deep learning architectures, integral to tasks such as classification, regression, and sequence modeling. Structurally, an MLP comprises fully connected layers where each neuron in a layer connects to every neuron in the subsequent layer. Originally developed as one of the first neural network models Haykin (1994), MLPs remain essential in contemporary architectures like transformers and vision-based networks Vaswani et al. (2017); Dosovitskiy et al. (2020) due to their capacity to model complex relationships and approximate intricate functions.
2.1 Role and Structure of MLPs
MLPs consist of sequences of linear transformations interspersed with non-linear activation functions, where the linear transformation can be represented as $\mathbf{y}=\mathbf{W}\mathbf{x}+\mathbf{b}$. Without the non-linear functions, consecutive linear transformations would collapse into a single transformation, limiting the model’s ability to learn non-linear relationships Goodfellow (2016). Popular activation functions in MLPs include ReLU, parametric ReLU (PReLU), GeLU, ELU, Sigmoid, and Tanh, with ReLU widely used for its computational efficiency and mitigation of the vanishing gradient problem Nair & Hinton (2010); Glorot et al. (2011).
2.2 MLPs in Modern Architectures
MLPs are pivotal in advanced architectures like transformers, which utilize MLP layers following self-attention mechanisms to enhance representational capacity Vaswani et al. (2017). In Vision Transformers (ViTs), MLPs process image patch embeddings, enabling spatial and feature learning from image data Dosovitskiy et al. (2020). Architectures like MLP-Mixers further showcase the versatility of MLPs by leveraging them for token and channel mixing, demonstrating that MLPs alone can achieve competitive performance in vision tasks Tolstikhin et al. (2021).
2.3 Importance of Non-Linear Activations
The inclusion of non-linear activation functions between linear layers is essential for MLPs to capture complex, non-linear mappings. Activation functions like ReLU introduce discontinuities that prevent the model from being purely linear, while bounded functions like Sigmoid and Tanh enable stability in certain applications despite their tendency to saturate Glorot et al. (2011). By breaking up linear layers, non-linearities enable MLPs to approximate complex functions effectively, a quality exploited in diverse applications.
2.4 Applications and Challenges
MLPs are critical in classification heads for converting high-dimensional representations to class scores in tasks such as image classification and sentiment analysis Russakovsky et al. (2015). In regression tasks, MLPs predict continuous values and are widely used in financial forecasting and scientific applications. Despite these strengths, the dense nature of MLPs contributes significantly to the computational overhead in models, spurring the need for optimization techniques like pruning, quantization, and activation modulation Fan et al. (2019); Hubara et al. (2017). Activation modulation, in particular, can simplify models by modulating non-linear activations toward linearity, thus reducing computational demands without heavily impacting model expressiveness.
@@ 3 Methodology
This section presents LayerCollapse , a novel regularization technique designed to adaptively compress the model by reducing its depth, possibly improving generalization in the process. In LayerCollapse , we employ compression-aware training. This approach enables the extraction of a shallower model directly, without the need for finetuning.
As illustrated in Figure 2, LayerCollapse involves regularizing the complexity of a layer by linearizing its activation function when training or fine-tuning. This linearization enables the merging of adjacent linear layers into a single layer, thus effectively reducing the model’s depth. The resulting model has lower storage, and computation demands without the need for software/hardware support.
To gain intuition on the effect that the non-linearities have on the model’s representation power and complexity, we present a toy scenario. Consider a 2-layer MLP with a hidden dimension of 500 and input and output dimensions of 1, employing a ReLU non-linearity. This network can be used to fit a polynomial function $f(x)=(x-3)(5x+2)(x+1)$ with samples having Gaussian noise. We apply gradient descent to fit the MLP to this function, aiming to uncover the underlying function $f(x)$. The results of this experiment are shown in Figure 1. In this demo, the regular model is described above. The $n\%$ linear model is a leaky ReLU with a negative slope of $n\%$. We observe that at $n=0.8$, the complexity of the 2-layer network is reduced such that it does not overfit to noise, closely approximating the underlying function. Conversely, at $n=0.98$, the reduced complexity of the model leads to underfitting. Setting $n=100\%$ results in a linear function, $Activation(x)=x$, effectively simplifying the 2-layer MLP to just a linear transformation since the product of the matrices $L_{1}^{1\times m}$ and $L_{2}^{m\times 1}$ results in a scalar. Here, $L_{1}\in\mathbb{R}^{1\times m}$ and $L_{2}\in\mathbb{R}^{m\times 1}$ are the weight matrices of the first and second layers.
Following the intuition from this experiment, we can see the effect of the negative slope on the model’s expressiveness. We design a regularizer that adaptively controls the negative slope of the MLP, preventing the model from overfitting to the noise. If for any layer we find that the negative slope is one, we can remove that layer from the network following the procedure in Algorithm 1 without losing any performance. The resulting network will have a shallower depth, fewer parameters, and lower resource requirements for execution.
3.1 Formal Definition
We define a Collapsible MLP as: Two consecutive fully connected layers utilizing a Parametric Rectified Linear Unit (PReLU) He et al. (2015) as the activation. The Parametric Rectified Linear Unit (PReLU) can be expressed as:
@@ 3.1 Formal Definition
where, $X$ is the input random variable and $Y$ is the output random variable. $W_{1}$ and $W_{2}$ are the weight matrices and $b_{1}$ and $b_{2}$ are the bias vectors for the first and second fully connected layers, respectively. $\alpha$ can be used to make the activation linear. We define $Y_{linear}$ as:
@@
We define the non-linearity error as the squared difference of $Y_{\alpha}$ and $Y_{linear}$. In Appendix A, we show that $\forall\delta\in[0,1]$
@@
where, $\sigma_{max}(W_{2}W_{1})$ is the largest singular value of $W_{2}W_{1}$ By choosing $\delta$ arbitrarily small, probability of error is bounded proportional to $(1-\alpha)^{2}$ with probability one. We incorporating $(1-\alpha)^{2}$ into our loss function to optimize for the non-linearity loss as well as the label cross-entropy loss. This loss term serves to align $Y_{\alpha}$ with $Y_{linear}$, compressing the model into a shallower model as demonstrated in Figure 2. In simpler terms, as $\alpha$ approaches one, substituting $Y_{linear}$ with $Y_{\alpha}$ incurs a bounded error probability, the error probability is proportional to $(1-\alpha)^{2}$ times a constant.
In practical applications, Dropout Srivastava et al. (2014) and BatchNormalization Ioffe & Szegedy (2015) are frequently used between the layers of a network. Dropout works by temporarily deactivating random neurons during training, which helps spread the learning across different neurons to avoid over-reliance on any particular subset of them. It is important to note that Dropout is not active during model deployment, meaning it behaves as if it is not there. This characteristic means that Dropout doesn’t affect the layer collapsing process and can be left out of consideration.
@@
where, $X$ and $Y$ are the input and output random variables. $\gamma$ and $\beta$ are learnable parameters, and $E[X]$ and $Var[X]$ are the mean and variance of the current batch.
Algorithm 1 outlines the specific steps to be taken for adjusting weights and biases when BatchNormalization is present.
3.2 Compression Ratio
Compression ratio for the collapsed layer can be calculated as:
$CR=1-\frac{n_{in}\times n_{out}}{n_{hidden}\times(n_{in}+n_{out})}$ (7)